QUESTION 9 Suppose there are three potential users of a freeway: Mr. 1. Mr. 2, and Mr. 3. The cost of the best alternative route for each commuter is as follows Commuter Alernate Cost Mr. 2 Mr. 3 S5 The average cost AC of using the freeway (i e, the cost per car) as a function of traffic volume T is as follows: T AC Mr. 1 $4 Mr 2 $4.3 Mr. 3 $9 Find the equilibrium allocation of trafflc between the freeway and alternate routes. Mr. 12 should stay on alternative route. should stay on freeway and Mr.3 Example if you thinlk Mr.l should stay on freeway and Mr. 2 & M:3 should stay on alternative ways, you should fill n the first blank: I: fill in the second blank 2,3 with a comma.) QUESTION 10 Followed by question S, Compute the total commuting cost for all commuters for the following four allocatious of traffic. Total cost is the cost incurred by freeway users plus the cost incurred by means no, please fill in the blanks) commuters who use their alternate routes (y means yes and n On freeway On alternate routes Total cost No one All Mr. 1 Mr. 2, Mr. 3 Mr. 1, Mr. 2 Mr. 3 All Social Optimal (fill in y or n) Page 267 No one
Answer #1

Case 1 : Total cost = 7 + 5 + 1 = $13.

Case 2: Total cost = 4 + 5 + 1 = $10.

Case 3: Total cost = 4 + 4.3 + 1= $9.3.

Case 4: Total cost = 4 + 4.3 + 9 = $17.3.

In case 3, total cost of the society is minimised at $9.3. so this is the social optimal case and rest are not.